# Getting Python Types Right with RightTyper

### v2

@@ 4.5. Threats to Validity
Our evaluation depends on the correctness of our TypeSim reimplementation and the ground truth used. We mitigate this by releasing our implementation and TypeEvalPy corrections for community review. The choice of sampling parameters (k=5k{=}5, $\lambda{=}2\,\mathrm{Hz}$, $\tau{=}0.05$) may affect results; we use defaults that balance overhead and coverage based on preliminary experiments.
@@
Our real-world evaluation covers two Python packages; while both are development tools, they exercise diverse language features and contain annotations refined through years of production use. Results may differ for other domains (e.g., web frameworks or ML code), though RightTyper’s language-agnostic design imposes no domain-specific limitations. Additionally, when execution is driven by test suites, RightTyper’s coverage depends on test quality: codebases with limited test coverage may yield incomplete annotations, while tests that exercise unintended behavior may result in annotations that reflect that behavior.

### v3

@@ 4.7. Threats to Validity
Our evaluation depends on the correctness of our TypeSim reimplementation and the ground truth used. We mitigate this by releasing our implementation and TypeEvalPy corrections for community review. The choice of sampling parameters (k=5k{=}5, $\lambda{=}2\,\mathrm{Hz}$, $\tau{=}0.05$ for event sampling; minimum 24 samples, budget cap 128 for container sampling) may affect results; we use defaults that balance overhead and coverage based on preliminary experiments.
@@
Our real-world evaluation covers five Python packages spanning development tools, libraries, and a web framework; they exercise diverse language features and contain annotations refined through years of production use. Results may differ for other domains (e.g., ML code), though RightTyper’s language-agnostic design imposes no domain-specific limitations. Additionally, when execution is driven by test suites, RightTyper’s coverage depends on test quality: codebases with limited test coverage may yield incomplete annotations, while tests that exercise unintended behavior may result in annotations that reflect that behavior.